Ticket Number: HEA-NOC/20080517-1 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: ESB
Ticket Opened: 20080517 09:41 UTC+1 Problem Start: 20080517 09:05 UTC+1
Ticket Closed: 20080518 19:21 UTC+1 Problem End: 20080518 11:08 UTC+1
Layer2 link from the Cork PoP to the HMRC went down at 09:05.
No traffic between HMRC and UCC and hence to the rest of the network.
20080517 09:38 BN Informed UCC. L2 equipment appears to be fine and HMRC
is connected on the Cork ring, so this suggests a power issue on-site.
Link will be monitored.
20080517 13:02 BN This was due to a power outage in the Cork area. The
connection will be monitored.
20080518 19:21 BN Connection has been stable and no further outages
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20080517-1
HEAnet Limited. Registered in Ireland, No. 275301.